# Emergence of Fluctuation Relations in UNO

### v1

@@ V Discussion
Our numerical experiment has found Crooks’ relation occurs in a card game, for a non-thermodynamical analog of work and for dissipation in a finite non-Markovian bath. First, in order to gain some intuition on why Crooks’ theorem holds in such a case, we compare our process to the familiar Markovian random walk (Section V.1). We then trace the origin of the observed ‘temperature’ (be it positive or negative) to the relation between the starting point of the process and the typical hand size (Section V.2). Finally, we discuss how the plus-cards affect the extent in which Crooks’ relation is adhered to (Section V.3).
Figure 4: A comparison of $\ln{\left(\frac{P_{F}}{P_{R}}\right)}$ ) against $W$ before and after ’Markovianisation’. The particular transition in the figure is $5\rightarrow 2$. For other transitions, we observed qualitatively identical behaviors.
(a) UNO under official rules.
(b) UNO under official rules, without +2 cards and +4 variation wildcards.
Figure 5: The slopes $\beta(x\rightarrow y)$ of the linear regression lines for different transitions $x\rightarrow y$, for $1\leq x,y\leq 10$. In the original version of Crooks’ fluctuation theorem, this slope is the inverse temperature $\beta$ of the thermal bath.
Figure 6: Histogram of Alice’s hand size throughout all of the games under the complete official rules. Upper graph: forward process; lower graph: reverse process. Having sampled each process for over $N=10^{6}$ runs, the error bars are $\sqrt{p(1-p)/N}\approx 10^{-4}$, too small to be seen at the scale of the plot.
@@ V.1 Comparison with a Markovian random walk
It is clear that, given our demarcations of the system, bath and designated state space, the process of playing UNO is non-Markovian. Indeed, the probability of losing/drawing a card on a given round depends not only on the number of cards at hand, but the actual cards themselves. Moreover, any given trajectory of an UNO game results in a memory that is embedded in the cards of the play deck, main deck and hands of other players. Transitions of the target system’s hand-size thus depend on this hidden memory, in a way that cannot be reduced to the previous, surface state, that it is transiting from.
Now, it may be insightful to quantify the extent of this non-Markovianity. To do this, we compare the original process to a ‘Markovianised’ version. The Markov transition matrix $M_{x,y}=P(y|x)$ will be given by the averaged transition probability of different hand sizes in a single round. We obtain these probabilities by averaging the probability of the $x\rightarrow y$ transition in a single round, regardless of the ‘actual’ state of the hand. What we obtain from this matrix is a Markov chain built from the UNO statistics where the probability of losing/drawing a card only depends on the current hand size. The Markov chain thus obtained is a finite, nonuniform random walk process. The nonuniformity can be understood through the fact that, in each round, Alice can only lose or draw one card, but the probabilities depend on the number of cards she currently has. For example, when she has 10 cards, it is more likely that she can play something compared to when she only has 2 cards. Thus, the transitions are dependent on position. It is finite since once she reaches state 0, the game ends.
@@
With this, we rerun the simulations as before, but for the Markov chain. A comparison of the Crooks fluctuation graphs obtained before and after Markovianisation can be seen in Fig. 4. It is clear that the plots behave similarly in the small $W$ regime but differ as $W$ increases.
@@
The analogy with a Markovian random walk may provide some intuition on the origin of the Crooks’ relation. Indeed, while a general treatment of a finite, nonuniform random walk is not trivial, it can be shown that Crooks’ theorem holds in an infinite, uniform random walk, when work corresponds to first passage time (see Appendix B).
With this, we see how the Crooks-like behavior emerges from the random-walk character in the intermediate $W$-regime. It is notable that this random-walk behaviour actually produces power law distributions in $W$ for each transition from $x$ to $y$ (3), as opposed to Gaussian distributions observed in physical fluctuation theorem experiments. This emphasizes that the settings which adhere to Crooks’ and Jarzynski’s theorems go beyond the typical thermodynamic case.
@@ V.2 The analog of temperature
In order to understand this behavior, we plot the histogram of our coarse-grained variable (Alice’s hand’s size) throughout all the games, for both the forward and the reverse process (Fig. 6). We observe that the most probable values are 3, 4 and 5. There is also a slight difference between the forward and reverse processes, the maximum being 3 for the former and 4 for the latter.
With this, we can explain that the most obvious structure in Fig. 5 as follows: $\beta(x\rightarrow y)$ is positive for transitions that move the hand size away from the most probable number $\mu\approx 3\sim 4$, i.e. if $|y-\mu|>|x-\mu|$; it is negative for transition that bring the hand size closer to it, i.e. if $|y-\mu|<|x-\mu|$. Also, $|\beta|$ is typically larger if the transition happens far away from $\mu$.
@@ V.3 The effects of plus-cards
As mentioned before, the removal of the plus-cards create some significant differences in the Crooks’ graph, which are: the extension of the support into both odd and even numbers, the oscillatory behaviour of the lines at small values of $W$, and the longer tails of the Crooks graphs. The first two are related and can be understood by our convention to count the action of drawing due to a +2 or +4 as a single deck interaction. Because of this, transitions that used to be possible in even (odd) steps can now be achieved in odd (even) steps too, given that Alice receives a plus-card. However, the probability of receiving a plus-card is relatively small, and most transitions still happen in the parity of $y-x$. This gives an oscillatory behaviour in both $P_{F}$ and $P_{R}$ between the two parity, and this behaviour carries over to the Crooks graphs. This behaviour, however, dies down at large values of $W$ as the chance of receiving a plus-card gets higher and higher the longer the trajectory is. Finally, the plus-cards also makes it harder to shed cards in general. Alice’s hand-size has a greater tendency to drift back and forth rather than decrease with regularity. This leads to generally longer paths and larger deck interactions.
@@
Overall, the addition of plus-cards leads the game away from a nice smooth behaviour in the Crooks graphs. This is expected as the plus-cards give a jumping behaviour to otherwise smooth trajectories of the hand size.
@@
We add that the same experiment was run for Crazy Eights (see Appendix C). Since it is played with standard playing cards without any analogue to plus-cards, the Crooks plots of Fig 7 have no significant oscillatory features, as expected.
VI Conclusion
@@
We have shown that Crooks’ fluctuation relation can be found in non-physics scenarios, even in a game of cards. Using the card-shedding game of UNO, we first replaced the thermodynamical notion of work with the number of interactions of one player with the deck. The bath, into which heat is gained from and dissipated into, is replaced by the cards held by the other players or sitting in the deck. Then we defined the forward process as a behavior or protocol of a player that is relatively successful at shedding cards, and the reverse process as the converse of that forward protocol. With these definitions, we run simulations of the game and collected the statistics of “work” for both the forward and the reverse process. These statistics, which obey a power law, eventually show that the definition of work that we adopted obeys Crooks’ fluctuation theorem for those processes.
@@
While a Crooks-like behavior is expected for a Markovian random walk, our process is non-Markovian, because the bath is finite and with memory. The unconventional nature of the bath is further demonstrated by the fact that the “temperature” (the inverse of the slope of the Crooks plot) can be either positive or negative, depending on the transition probabilities of the hand size. This behavior was explained as an artefact of the most common state being around the hand-size of $3$ and $4$.
@@
Our work establishes another bridge connecting game theory and thermodynamics [33, 34, 35, 36], and provides a new example of the relevance of tools from stochastic thermodynamics outside thermal physics [37, 38]. Going forward, we notice that here we fixed a simple strategy for the player and focused on establishing an experiment in analogy with thermodynamical physics. It would be interesting to see a deeper connection between different strategies of the player and their effects on the fluctuation relation. Moreover, a more extensive survey of different games would be helpful in elucidating the significance and generality of Crooks’ fluctuation theorem.
The codes used to run the simulations can be found online  [^3] https://github.com/PeterSidajaya/uno-fluctuation.
@@ Acknowledgments
We thank Alexia Auffèves, Francesco Buscemi, Andrés Ducuara, Hyukjoon Kwon, Dario Poletti and Antonios Varvitsiotis for helpful discussions and suggestions.
This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; and by the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). This project started in the context of the Science Research Programme (SRP), a collaboration between the NUS Faculty of Science, NUS College of Design and Engineering, and the Gifted Education Branch of the Ministry of Education, Singapore. We also thank Chan Xin Yu for her support in this collaboration.
@@ Appendix A Strategies used
In order to effect the forward and reverse processes, we created two algorithms, one that aims to win (forward) and another that aims to lose (reverse). Note that in UNO, it is typically legal to draw a card even when you can play a card from your hand. Thus, a possible reverse algorithm is just to draw a card every round. However, to make the reverse process less trivial we will force Alice to play when there is a possible legal move in her hand.
One of the strategy to win UNO is to keep any wildcard on the hand for as long as possible. This is because of the card’s flexibility to be played regardless of the state of the gamepile. Moreover, when playing a wildcard one should generally choose a colour that is most numerous in one’s hand. This is because colour changes is infrequent and colour tends to stay the same between subsequent rounds. These two observations constitute the strategy that Alice uses for the forward process. For the reverse strategy, Alice tries to throw away any wildcard she has as soon as possible. Moreover, she then chooses the least numerous colour on her hand, in the hope of being forced to draw a card as soon as possible.
The two algorithms are shown in Algorithm 1 and 2. Against three other bots playing completely randomly the win rate of the forward and reverse processes for the official UNO game are 33.8% and 24.7%, respectively.
@@
                   $X\leftarrow$ most common colour on hand;
                   play the wildcard and declare $X$ as the next colour;
@@
Algorithm 1 The forward process
@@
            $X\leftarrow$ least common colour on hand;
             play the wildcard and declare $X$ as the next colour;
@@
Algorithm 2 The reverse process
@@
Appendix B Crooks’ relation for the time of first passage in an infinite random walk
@@
Consider a random walk process on the lattice of integer numbers =$\mathds{Z}=\{...,-2,-1,0,1,2\}$. On each timestep, the probability of the walker going to the right ($+1$) is $p$, while the probability of it going to the left ($-1$) is $q=1-p$. In analogy with the definition of work adopted in the main text for the UNO game, let us define work as the first hitting/passage time from the initial to the final state, and denote this by the random variable $N$. Given an initial point $i$ in the lattice and given that the walker has moved from $i$ to $i+1$, what we want to find is the distribution of $P(N=2n-1|i\rightarrow i+1)$. By the method of generating function, it can be derived that [40]
$P(N=2n-1|i\rightarrow i+1)=\frac{(2n-3)!!}{n!}2^{n-1}p^{n}q^{n-1},$ (3)
where !!double-factorial!!!! is the double factorial, the product of all positive integers with the same parity as $n$.
@@
So far this only concerns one process. To make a fluctuation relation, let us define two random walks on two different lattices, say $A$ and $B$, with different transition probabilities, say $p_{1},q_{1}$ and $p_{2},q_{2}$. Taking the logarithm of the ratio of the first passage times, we would quickly find out
$\ln\left(\frac{P_{A}(N=2n-1|i\rightarrow i+1)}{P_{B}(N=2n-1|i\rightarrow i+1)}\right)$
$=n\ln\left(\frac{p_{1}q_{1}}{p_{2}q_{2}}\right)-\ln\left(\frac{q_{1}}{q_{2}}\right)$
Taking analogies with Crooks, we can define
$\beta=\ln\left(\frac{p_{1}(1-p_{1})}{p_{2}(1-p_{2})}\right),\quad\Delta F=\frac{1}{\beta}\ln\left(\frac{(1-p_{1})}{(1-p_{2})}\right).$ (4)
As of this point, we have not yet elaborated on the relations of the two processes. To get a more meaningful interpretation, the second random walk should be the ‘reverse’ of the first one. For example, let $G(p,q)$ be the graph on which the forward process is done. Define $P_{F}=P(N|i\rightarrow i+1;G(p_{1},q_{1}))$ as the forward process. Next, define the reverse as $P_{R}=P(N|i+1\rightarrow i;G(p_{2},q_{2}))$. This is simply the first passage time of the walker going back to the initial state from the end state with a different set of transition probabilities. Note that
$P(N|i+1\rightarrow i;G(p_{2},q_{2}))=P(N|i\rightarrow i+1;G(q_{2},p_{2})).$
Now, if the forward and reverse processes are done on the same lattice, we would have $p_{2}:=p_{1}$ and $q_{2}:=q_{1}$. In this particular way of reversing, with Eq. (4) we will get $\beta=0$. In general, without context, there is no unique way to reverse a random walk process. However, as long as the reverse process is still a random walk, the linear Crooks-like relation of Eq. (4) will hold.
Appendix C Crazy Eights
(a) Crooks plot for Crazy Eights.

### v2

@@ V Discussion
Our numerical experiment has found Crooks’ relation occurs in a card game, for a non-thermodynamical analog of work and for dissipation in a finite non-Markovian bath. First, in order to gain some intuition on why Crooks’ theorem holds in such a case, we compare our process to the familiar Markovian random walk (Sec. V.1). We then trace the origin of the observed “temperature” (be it positive or negative) to the relation between the starting point of the process and the typical hand size (Sec. V.2). Finally, we discuss how the plus-cards affect the extent in which Crooks’ relation is adhered to (Sec. V.3).
Figure 4: The probability mass functions (pmf) of different transitions for the forward (top) and reverse (bottom) processes. Similar to a random walk (see Appendix C), the pmf is in the form of an exponential distribution. Here, only the transitions in the full game with the plus-cards are shown.
Figure 5: The colormap transition probabilities of the “Markovianized” UNO for the forward process games with plus-cards. The two brightest diagonal lines correspond to the transition probabilities of playing one card or drawing one card. The other two diagonals correspond to the transition probabilities of drawing multiple cards at once due to the plus-cards.
Figure 6: A comparison of $\ln{\left(\frac{P_{F}}{P_{R}}\right)}$ ) against $W$ before and after “Markovianization”. The particular transition in the figure is $5\rightarrow 2$. For other transitions, we observed qualitatively identical behaviors.
@@ V.1 Comparison with a Markovian random walk
Figure 7: The slopes $\beta(x\rightarrow y)$ of the linear regression lines for different transitions $x\rightarrow y$, for $1\leq x,y\leq 10$ for (a) UNO under official rules and (b) UNO without the +2 and +4 cards. In the original version of Crooks’ fluctuation theorem, this slope is the inverse temperature $\beta$ of the thermal bath.
If we had defined our state space to be the full state of the game without any coarse graining, the whole process of playing UNO would be Markovian. However, because of our coarse graining process, the process is now non-Markovian. The probability of a given trajectory depends on a memory that is embedded in the cards of the play deck, main deck and hands of other players, which are not taken into account in the state space.
To make further analysis easier, it may be insightful to see whether the non-Markovian effect is negligible. Thus, we compare the original UNO process to a ‘Markovianised’ approximation. The Markov transition matrix $M_{x,y}=P(y|x)$ is obtained from the data by averaging the transition probability between different hand sizes $x\longrightarrow y$ in a single round. In actual play this probability depends variables other than Alice’s current hand size, such as the type of cards in her hand, but the ’Markovianised’ approximation effectively ignores these other variables. What we obtain from this matrix is a Markov chain built from the UNO statistics where the probability of losing/drawing a card only depends on Alice’s current hand size. The transition probabilities can be seen in Fig. 5
The Markov chain thus obtained is a finite, nonuniform random walk process (with smaller “jump” transitions allowed by the plus-cards) with an absorbing boundary at 0. As seen from gradient of color in the transition matrix of Fig. 5, the probability of playing a card in a given round is monotonically increasing with the number of cards on hand. This is intuitive as it is more likely that Alice is able to play something when she has 10 cards compared to when she only has three. Finally, it is finite since once she reaches state 0, which acts as an absorbing state, the game ends.
Our system bears resemblance to the continuous time and space Ornstein-Uhlenbeck process, which can also give an exponential first passage times under certain conditions [36]. However, fully analytical treatment of a stochastic process such as these is non-trivial and lies outside the scope of the paper. For a more complete treatment of first passage times in stochastic processes, we refer the readers to references such as [37]
@@
With this, we rerun the simulations as before, but for the Markov chain. A comparison of the Crooks’ fluctuation graphs obtained before and after Markovianisation can be seen in Fig. 6. The values are similar for small values of $W$, then differ for larger values. Thus, we see how the Crooks-like behavior emerges from the random-walk character in the intermediate $W$ regime.
@@
The analogy with a Markovian random walk may provide some intuition on the origin of the Crooks’ relation. Indeed, although a general treatment of a finite, nonuniform random walk is not trivial, it can be shown that Crooks’ theorem holds in an infinite, uniform random walk, when work corresponds to first passage time (see Appendix C). It is notable that this random-walk behavior actually produces exponential distributions in $W$ for each transition from $x$ to $y$ (see Fig. 4), as opposed to Gaussian distributions observed in physical fluctuation theorem experiments. This emphasizes that the settings that adhere to Crooks’ and Jarzynski’s theorems go beyond the typical thermodynamic case.
@@ V.2 The analog of temperature
In order to understand this behavior, we plot the histogram of our coarse-grained variable (Alice’s hand size) throughout all the games, for both the forward and the reverse process (Fig. 8). We observe that the most probable values are 3, 4 and 5. There is also a slight difference between the forward and reverse processes, the maximum being 3 for the former and 4 for the latter.
With this, we can observe that the basic structure in Fig. 7 is as follows: $\beta(x\rightarrow y)$ is positive for transitions that move the hand size away from the most probable number $\mu\approx 3\sim 4$, i.e., if $|y-\mu|>|x-\mu|$; it is negative for transitions that bring the hand size closer to it, i.e., if $|y-\mu|<|x-\mu|$. Also, $|\beta|$ is typically larger if the transition happens far away from $\mu$.
An interesting point to note is that, in the case of an infinite uniform random walk (see Appendix C), the slope of the fluctuation relations is the same regardless of the transition, as one might expect from a physical temperature. In turn, this shows why a single “temperature” is not expected in UNO: even when Markovianised (Sec. V.1), the transition probabilities are not uniform but state dependent; and the walk is bounded on one side. We leave it for further work to explore whether the analogy with temperature, natural in the fluctuation theorem analysis, can be fruitful beyond it.
Figure 8: Histogram of Alice’s hand size throughout all of the games under the complete official rules. (Top) Forward process; (bottom) reverse process. Having sampled each process for over $N=10^{6}$ runs, the error bars are $\sqrt{p(1-p)/N}\approx 10^{-4}$, too small to be seen at the scale of the plot.
@@ V.3 The effects of plus-cards
As mentioned before, the removal of the plus-cards create some significant differences in the Crooks’ graph, which are: the extension of the support into both odd and even numbers, the oscillatory behavior of the lines at small values of $W$, and the longer tails of the Crooks graphs.
@@
The first two are related and can be understood by our convention to count the action of drawing due to a +2 or +4 as a single deck interaction. Because of this, transitions that used to be possible in even (odd) steps can now be achieved in odd (even) steps too—given that Alice receives a plus-card. However, the probability of receiving a plus-card is relatively small, and most transitions still happen in the parity of $y-x$. This gives an oscillatory behavior in both $P_{F}$ and $P_{R}$ between the two parity, and this behavior carries over to the Crooks graphs. This behavior, however, dies down at large values of $W$ as the chance of receiving a plus-card gets higher and higher the longer the trajectory is. This initial zigzag behavior can be seen more clearly in the pmf of $W$ in each of the processes in Fig. 4. Finally, the plus-cards also make it harder to shed cards regularly. Alice’s hand size has a greater tendency to drift back and forth rather than decrease with regularity. This leads to generally longer paths and larger deck interactions.
@@
Overall, the addition of plus-cards leads the game away from a smooth behavior in the Crooks’ graphs. This is expected as the plus-cards give a stepwise behavior to otherwise smooth trajectories of the hand size.
@@
We add that the same experiment was run for Crazy Eights (see Appendix D). Since it is played with standard playing cards without any analog to plus-cards, the Crooks plots of Fig 10 have no significant oscillatory features, as expected.
VI Conclusions
@@
We have shown that Crooks’ fluctuation relation can be found in non-physics scenarios, even in a game of cards. Using the card-shedding game of UNO, we first replaced the thermodynamical notion of work with the number of interactions of one player with the deck. The bath, into which heat is gained from and dissipated into, is replaced by the cards held by the other players or sitting in the deck. Then, we defined the forward process as a behavior or protocol of a player that is relatively successful at shedding cards, and the reverse process as the converse of that forward protocol. With these definitions, we run simulations of the game and collected the statistics of “work” for both the forward and the reverse process. These statistics, which obey an exponential distribution show that the definition of “work” obeys Crooks’ fluctuation theorem in these processes.
@@
While a Crooks-like behavior is expected for a Markovian random walk, our process is non-Markovian, because the bath is finite and with memory. The unconventional nature of the bath is further demonstrated by the fact that the “temperature” (the inverse of the slope of the Crooks plot) can be either positive or negative, depending on the transition probabilities of the hand size. This behavior was explained as a result of the most common state being around the handsize of $3$ and $4$.
Our study establishes another bridge connecting game theory and thermodynamics [38, 39, 40, 41, 42, 20, 43, 44], and provides an example of the relevance of tools from stochastic thermodynamics outside thermal physics [45, 46]. In particular, our study shows how such comparisons across game theory and physics produce questions that can lead to new insights—in this case, connections between random walks, their boundedness, Markovianity, Crooks’ fluctuation theorem, and temperature.
@@
Going forward, we notice that here we fixed a simple strategy for the player and focused on establishing an experiment in analogy with thermodynamical physics. It would be interesting to see a deeper connection between different strategies of the player and their effects on the fluctuation relation. Moreover, a more extensive survey of different games, or even beyond games, would be helpful in elucidating the significance and generality of Crooks’ fluctuation theorem.
@@ Acknowledgments
We thank Alexia Auffèves, Francesco Buscemi, Andrés Ducuara, Géraldine Haack, Hyukjoon Kwon, Dario Poletti, Paul Skrzypczyk and Antonios Varvitsiotis for helpful discussions and suggestions.
This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; by the same National Research Foundation, Singapore, through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009); and by the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002).
This project started in the context of the Science Research Programme (SRP), a collaboration between the NUS Faculty of Science, NUS College of Design and Engineering, and the Gifted Education Branch of the Ministry of Education, Singapore. We also thank Chan Xin Yu for her support in this collaboration.
Code and Data Availability
The codes used to run the simulations and generate the data can be found online  [^3] https://github.com/PeterSidajaya/uno-fluctuation. Requests and correspondences regarding the code should be addressed to P.S.
@@ Appendix A Strategies used
In order to effect the forward and reverse processes, we created two algorithms, one that aims to win (forward) and another that aims to lose (reverse). As noted earlier, it is actually legal to draw a card even when you can play a card from your hand. Thus, a possible reverse algorithm is to draw a card every round. However, to make the reverse process less trivial we will force Alice to play whenever there is a possible legal move.
One of the strategies to win UNO is to keep any wildcard on the hand for as long as possible. This is because the card can be played regardless of the state of the gamepile. Moreover, when playing a wildcard one should generally choose a color that is most numerous in one’s hand. This is because color rarely changes and tends to persists between subsequent rounds. These two observations constitute the strategy that Alice uses for the forward process. For the reverse strategy, Alice tries to throw away any wildcard she has as soon as possible. Moreover, she then chooses the least numerous color on her hand, in the hope of being forced to draw a card as soon as possible.
The two algorithms are shown in Algorithm 1 and 2. Against three other bots playing completely randomly the win rate of the forward and reverse processes for the official UNO game are 31.7% and 24.7%, respectively.
@@
                   $X\leftarrow$ most common color on hand;
                   play the wildcard and declare $X$ as the next color;
@@
Algorithm 1 The forward process.
@@
            $X\leftarrow$ least common color on hand;
             play the wildcard and declare $X$ as the next color;
@@
Algorithm 2 The reverse process.
Appendix B Different kinds of bath
Figure 9: The Crooks’ plot for different versions of the bath (a) UNO game with other players playing smartly (compare with Fig. 3(a)) (b) UNO game with two players and (c) UNO game with six players. Each of the data for these variants was obtained from two million runs.
In the main text, we only considered four-player games where the other players always play randomly. Here, we consider two different variations of the bath (the other players): one with smart players and another with a different number of players.
Previously, we only considered the case where the other players other than Alice always played randomly without any strategy. One might then ask the question of whether the randomness of the other players ‘adds’ to the stochasticity of the bath and affects our result. The short answer to this question is: No. Figure 9(a) shows the Crooks’ plot obtained with the same method as in the main text but with all the other players using Algorithm 1 instead of playing randomly. As can be seen, there is little to no noticeable difference with Fig. 3(a).
While we only considered four-player games previously, UNO is actually designed for two to ten players, thus we also investigated what happens when we change the number of players. The Crooks’ plot for a game of UNO with two and six players can be seen in Figs. 9(c) and  9(b), respectively. Two trends can be seen here: With more cards, the length of the graph shortens and the gradients for each transition go down. The difference in length can be understood from the fact that, in games with more players, the trajectories of Alice tend to be shorter. That is, in games with more players, Alice plays less in each game.
@@ Appendix B Different kinds of bath
Appendix C Crooks’ relation for the time of first passage in an infinite random walk
@@
Let us consider a random walk process on the lattice of integer numbers =$\mathds{Z}=\{...,-2,-1,0,1,2\}$ where transitions exist only between neighboring numbers. In each timestep, the probability of the walker going to the right ($+1$) is $p$, while the probability of it going to the left ($-1$) is $q=1-p$. In a Markov chain, the first passage/hitting time $N$ is the number of steps taken by a walker to first reach $f$ from a starting point $i$. What we are concerned with is the distribution of $P(N|i\rightarrow i+k)$, i.e., the distribution of the first passage time given that the walker has moved from $i$ to $i+k$. The closed form expression is given by [48]
$P(N=n|i\rightarrow i+k)=\frac{|k|}{n}\binom{n}{\frac{n+k}{2}}p^{\frac{n+k}{2}}q^{\frac{n+k}{2}-k},$ (3)
and it is zero when $\frac{n+k}{2}$ is not a nonnegative integer. This distribution is an exponential distribution (compare with in Fig. 4).
@@
To make a fluctuation relation, we need two random walks $A$ and $B$ with different transition probabilities, $(p_{1},q_{1})$ and $(p_{2},q_{2})$. Taking the logarithm of the ratio of the first passage times, we get
$\ln\left(\frac{P_{A}(N=2n-1|i\rightarrow i+k)}{P_{B}(N=2n-1|i\rightarrow i+k)}\right)$
$=\frac{n+k}{2}\ln\left(\frac{p_{1}q_{1}}{p_{2}q_{2}}\right)-k\ln\left(\frac{q_{1}}{q_{2}}\right)\,;$
where the analogy with Crooks with $n\equiv W$ yields
$\beta=\frac{n}{2}\ln\left(\frac{p_{1}(1-p_{1})}{p_{2}(1-p_{2})}\right),\quad\beta\Delta F=-\frac{k}{2}\ln\left(\frac{p_{1}(1-p_{1})}{p_{2}(1-p_{2})}\right)+k\ln\left(\frac{(1-p_{1})}{(1-p_{2})}\right).$ (4)
In particular, $\beta$ is independent of $k$, i.e., it is the same for all transitions. We have mentioned this fact in Sec. V.2, and explained there why this is naturally not the case for UNO.
Finally note that, at this stage, we have not yet elaborated on the relations of the two processes. As is the case in general [18, 19] (and we have explained at the end of Sec. III why our experiment does not fall into this triviality trap), the detailed fluctuation theorem follows formally from studying $\ln(P_{A}/P_{B})$, without the need to provide any further connection between the two processes. Here, for the specific relation, we added the request that both processes be random walks.
For deeper interpretations, one may want to require further that process $B$ would be a reverse of $A$. There is no general way to reverse an abstractly defined random walk. Physically, one could think of a walk on an inclined plane (a native $p\neq 1-p$) with an additional drive, pushing in opposite directions for processes A and B (to get the effective $p_{1}\neq p_{2}$). We don’t need to pursue this analogy here.
Appendix D Crazy Eights
(a) Crooks’ plot for Crazy Eights.